# Cell Maps Representation FOR LUNG Adenocarcinoma GROWTH PATTERNS CLASSIFICATION IN WHOLE SLIDE IMAGES

### v1

@@ 1 Introduction
  • We introduce the concept of cell maps for predicting pathology specific tasks that can add generalizability to current machine learning approaches.
@@
Fig. 2: An overview of the proposed Model. First the WSI is divided into tiles, which are then processed using HoverNet to obtain cell locations and types, then a cell map is generated by stacking the masks of three cell types (neoplastic, non-neoplastic, connective). The generated cell map is used as an input to a Convolutional neural network (CNN) for growth pattern classification. * Colors of the cell masks are inverted for illustration purposes
@@ 2.2 Growth pattern classification
We finetuned a ResNet50, pretrained on ImageNet, after adapting the last fully connected layer to predict one of the six classes (lepidic, acinar, papillary, micropapillary, solid, and normal). Cell map tiles of size $256\times 256256 × 256$ were passed to the network, after applying random horizontal and vertical flipping. The model was trained for 20 epochs with a learning rate of 1e-5, which due to the simplicity of the input image representation, we found to be sufficient for the model to converge.
@@ 3.1 Dataset
Model AUCROC accuracy F1-macro AUCROC accuracy F1-macro
@@ 3.2 Results
    ResNet50: (the architecture used in [8] and [16]) pre-trained on ImageNet fed with H&E tiles at $20\times$20 × magnification. This architecture was used in
@@ 3.2.1 Weak validation
For our first experiment, we adopted a tile-based data splitting approach, where we mixed all the tiles and used 5-fold cross validation to prove the effectiveness of the proposed model and its comparability to the currently published work. Quantitative results are listed in Table1.
@@ 3.2.2 Strong validation
For our main experiments, we adopt a WSI-based data splitting approach, where we trained and evaluated each model 5 times. In each trial 6 WSIs were randomly sampled from the dataset to be the unseen test set, ensuring that it includes samples from all the six classes. The remaining slides were divided into tiles and split into 90% training and 10% validation set. The average and standard deviation of the performance measures (AUC-ROC, accuracy, and macro F1-score) are reported in Table 1.
It has been shown in Table 1 that when performing proper validation by adopting a WSI-based data splitting, the proposed cell maps approach performs better on unseen test sets compared to other methods, some of which perform worse than random guessing. Although using the H&E image directly for classification initially suggests better performance, when properly validating the models via a WSI-based splitting, the performance dramatically drops by over 60%. Conversely, when employing cell maps, the decline in performance is notably less, dropping by under 20%.
@@ 4 CONCLUSIONs
We proposed a new approach for LUAD growth pattern classification, that outperforms previously proposed approaches in the literature and baseline methods, when evaluated using WSI-based splitting. Additionally, we present a new representation of WSIs, called cell maps, which effectively captures cellular composition in more compact and lighter images. This innovative approach can significantly expedite the training of machine learning algorithms for some pathological tasks. Our future directions include acquiring more annotations to expand our dataset along with testing the model on external cohorts. We aim to project the model predictions on the entire WSI for prognosis and survival analysis.

### v2

@@ 1 Introduction
  • We introduce the concept of cell maps for predicting pathology specific tasks that can add generalizability to current machine learning approaches. Up to our knowledge, this is the first study that uses cell maps representation for tissue classification.
@@
Fig. 2: An overview of the proposed Model. First the WSI is divided into tiles, which are then processed using HoverNet to obtain cell locations and types, then a cell map is generated by stacking the masks of three cell types (neoplastic, non-neoplastic, connective). The generated cell map is used as an input to a Convolutional neural network (ResNet50) for growth pattern classification. * Colors of the cell masks are inverted for illustration purposes
@@ 2.2 Growth pattern classification
We finetuned a ResNet50, pretrained on ImageNet, after adapting the last fully connected layer to predict one of the six classes (lepidic, acinar, papillary, micropapillary, solid, and normal). Cell map tiles of size $256\times 256256 × 256$ were passed to the network, after applying random horizontal and vertical flipping. Training was optimized using an Adam optimizer and a cross entropy loss function. The model was trained for 20 epochs with a learning rate of 1e-5, which due to the simplicity of the input image representation, we found to be sufficient for the model to converge.
@@ 3.1 Dataset
Model AUCROC F1-macro accuracy AUCROC F1-macro accuracy
@@ 3.2 Results
    ResNet50: (the architecture used in [8] and [16]) pre-trained on ImageNet fed with H&E tiles at $20\times$20 × magnification.
@@ 3.2.1 Weak validation
This validation approach used in current literature involves employing a tile-based data splitting approach, where 5-fold cross validation is applied after all the tiles from different WSIs are mixed. We include the results of this weak validation solely for comparison purposes. We suggest strong validation should be used for testing robustness of the algorithms. All models was trained for 20 epochs with a learning rate of 1e-5.Training was optimized using an Adam optimizer and a cross entropy loss function. Quantitative results are listed in Table1.
@@ 3.2.2 Strong validation
For our main experiments, we adopt a WSI-based data splitting approach, where we trained and evaluated each model 5 times. In each trial 6 WSIs were randomly sampled from the dataset to be the unseen test set, ensuring that it includes samples from all the six classes. The remaining slides were divided into tiles and split into 90% training and 10% validation set. Training was optimized using an Adam optimizer and a cross entropy loss function. All models was trained for 20 epochs with a learning rate of 1e-5. The average and standard deviation of the performance measures (AUC-ROC, accuracy, and macro F1-score) are reported in Table 1.
It has been shown in Table 1 that when performing proper validation by adopting a WSI-based data splitting, the proposed cell maps approach performs better on unseen test sets compared to other methods, some of which perform worse than random guessing. Although using the H&E image directly for classification initially suggests better performance, when properly validating the models via a WSI-based splitting, the performance dramatically drops by over 60%. Conversely, when employing cell maps, the decline in performance is notably less, dropping by under 20%. One could argue that cell maps enable CNNs to concentrate on cell morphology without becoming overly fixated on tissue-specific characteristics.
The proposed cell maps approach was best at distinguishing non-tumor and solid pattern with an average f1-score of 0.98 and 0.97 respectively. On the other hand, the papillary pattern exhibited the highest misclassification error. This could be attributed to the fact that the samples originated from only four WSIs, and given the high heterogeneity of this pattern, the model might not have been exposed to a sufficient number of samples during training. We plan to address this issue by using a larger dataset in the extended version of this work.
@@ 4 CONCLUSIONs
We proposed a new approach for LUAD growth pattern classification, that outperforms previously proposed approaches in the literature and baseline methods, when evaluated using WSI-based splitting. Additionally, we present a new representation of WSIs, called cell maps, which effectively captures cellular composition in more compact and lighter images. This innovative approach can significantly expedite the training of machine learning algorithms for some pathological tasks. We have also shown that the current validation technique used in the literature to evaluate growth pattern classification models is weak and should be avoided; as it can generate misleading results. Our future directions include acquiring more annotations to expand our dataset along with testing the model on external cohorts. We aim to project the model predictions on the entire WSI for prognosis and survival analysis.